what is the value of cos(180-2theta)??
Answer by stanbon(75887)

what is the value of cos(180-2theta)
------
= cos(180)(cos(2t)) - sin(180)(sin(2t)
------------
= -cos(2t) - 0*sin(2t)
-----
= -cos(2t)
--------------
Cheers,
Stan H.
---------------